OpenStudy (anonymous):

John is x years old.Jack is 4 years younger than John.The product of the ages of John and Jack is 320.Form a quadratic equation. @sammixboo

sammixboo (sammixboo):

What did you get?

sammixboo (sammixboo):

Sorry I might take a minute or two to respond because I am having internet problems

OpenStudy (anonymous):

alrite :)

OpenStudy (anonymous):

idk :(

sammixboo (sammixboo):

I am confusing myself now. I know that John is 20 and Jack is 16, but I can't form the equation. @ganeshie8 can you help?

sammixboo (sammixboo):

@campbell_st

OpenStudy (campbell_st):

Jack = x -4. John = x The product x(x -4)= 320 Just distribute and then subtract 320 from both sides of the equation

OpenStudy (anonymous):

x^2-4x-320=0

OpenStudy (anonymous):

thnx @campbell_st @sammixboo

OpenStudy (campbell_st):

That's it... Reading the question there in no need to solve